I am working with a non-right triangle with Angle A and Angle B unknown. Angle C measures 36.5 degrees. Side a=13, b=?, c=7.8. The question states that sin(36.5)=0.5948, and cos(36.5)=0.8039. I am having trouble figuring out how to calculate sin A. Can someone out there please help me??? I have tried using the law of cosines but I ran into a dead end quickly on that try. Thank you for any help that you can provide!!
Answer by stanbon(48558)

I am working with a non-right triangle with Angle A and Angle B unknown. Angle C measures 36.5 degrees. Side a=13, b=?, c=7.8. The question states that sin(36.5)=0.5948, and cos(36.5)=0.8039. I am having trouble figuring out how to calculate sin A.
------------------------
Use the Law of Sines:
SinA /a = sinC/c
---------------
Sin(A)/13 = sin(36.5)/7.8
sin(A) = 13[sin(36.5)/7.8]
sin(A) = 0.991
===================
Cheers,
Stan H.